# Words with factor complexity $2n+1$ and minimal critical exponent

### v2

@@ 5 Proof of the Structure Theorem
We prove the Structure Theorem via an intricate case analysis. Each case is determined by a set $S\subseteq\Sigma_{3}^{*}$ where we assume that $S\subseteq F$, the set of factors omitted by $\mathbf{w}$}. As remarked in the previous section, it suffices to resolve the subcases $S=\{01,12,20,00\}$ and $S=\{11,22\}$.
In a given case, labelled by a set $S$, we proceed in one of three ways. We either:
@@
The concrete recipe for carrying out 1, 2, and 3 above is as follows:
@@
    Perform a backtrack search, looking for a low word $w$ of length 250 with no factors in $S$; if no such word exists, the case cannot arise.
@@
      If no such $n$ is found: We look for a morphism satisfying the conditions of the Morphism Lemma or the Dual Morphism Lemma. To do this, we find factors $b$ of $w$, with $|b|$ up to length 3, such that $b$ has at most 3 recurrent return words in $\mathbf{w}$}. We form a morphism from these 3 return words, and consider it and its conjugates.
      If one of the considered morphisms satisfies the conditions of the Morphism Lemma or the Dual Morphism Lemma the case is resolved.
    2. b.
@@
        If there are more than $2n+1$ needed factors then the condition on the complexity of $\mathbf{w}$} is violated and the case cannot arise.
@@
A python implementation of the above scheme in SageMath 9.3 (on a Microsoft Surface with Intel(R) Core(TM) i5-1035G4 CPU at 1.10GHz) resolved the case $S=\{11,22\}$ in about 15 seconds, and the case $S=\{01,12,20,00\}$ in about 4 seconds. ∎
@@
In fact, the implementation mentioned above kept track of sets $S$ and the reasoning at each of its steps, and produced ‘human readable’ proofs. We give the reasoning that was output for the case $S=\{01,12,20,00\}$ in an appendix. Further, the implementation kept track of all morphisms $h$ used. It may be the case that for some of these morphisms, $h({\mathbf{G}})$ is an eligible word with critical exponent $\mu$.
@@ 6 Proofs of Lemmas
The critical exponent of $\mathbf{u}$} is irrational, and therefore is not realized by any of its finite factors. Thus for any $\epsilon_{2}>0$ and any positive integer $N$, word $\mathbf{u}$} has a factor $uuu^{\prime}$ such that $u^{\prime}$ is a prefix of $u$ with $|u^{\prime}|>N$ and
@@
We use two preliminary lemmas in the proof of the Morphism Lemma (Lemma 4):
@@
Suppose that ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ is low. Let $a,b,c\in\Sigma_{3}^{+}$, and suppose that a final segment of $\mathbf{w}$} has the form $g({\mathbf{u}})$, some ${\mathbf{u}}\in\Sigma_{3}^{\omega}$, where $g=[a,b,c]$. Suppose further that:
@@
    Word $b$ is a prefix of $a$, which is a prefix of $c$.
@@
    Words $b$ and $c$ have a common suffix $s$ such that $|s|\geq|b|/2$.
Then $\mathbf{u}$} doesn’t contain factors $00$, $11$, 2121, or $22$.
@@
We treat the factors one at a time. For each factor we rule out 1-letter right extensions of that factor:
Factor $11$: If $11$ is a factor of $\mathbf{u}$}, then so is a factor $v$, where $v=110$, $v=111$, or $v=112$. In each case, $g(v)$ has the prefix bbbbbb which is a $3$ power. Since $\mathbf{w}$} doesn’t contain $3$ powers, $11$ doesn’t occur in $\mathbf{u}$}.
Factor $00$: If $00$ is a factor of $\mathbf{u}$}, then so is a factor $v$, where $v=000$, $v=001$, or $v=002$. In each case $g(v)$ has the prefix aabaab since $b$ is a prefix of $a$ and $c$. However, aabaab is an $r$ power, $r\geq 5/2$, since $b$ is a prefix of $a$ and $|b|\geq|c|/2\geq|a|/2$.
Factor $21$: If $21$ is a factor of $\mathbf{u}$}, then so is a factor $v$, where $v=211$, $v=210$, or $v=212$. In each case, $g(v)$ has the prefix cbbcbb. This has the suffix sbbsbb, which is an $r$ power, $r\geq 5/2$.
Factor $22$: If $22$ is a factor of $\mathbf{u}$}, then so is a factor $v$, where $v=220$, $v=221$, or $v=222$. In each case, $g(v)$ has the prefix ccbccb, which is an $r$ power, $r\geq 5/2$, since $b$ is a prefix of $c$ and $|b|\geq|c|/2$. ∎
It is convenient to work with $\gamma^{2}=[012,02,0102]$ rather than simply $\gamma$. If $g=[a,b,c]$, then $g\circ\gamma^{2}=[abc,ac,abac]$, and each image of a letter under $g\circ\gamma^{2}$ ends in $c$. This allows us to define $g^{\prime}=c(g\circ\gamma^{2})c^{-1}=[cab,ca,caba]$.
@@
Suppose that ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ is low. Suppose that $\mathbf{w}$} has the form $g({\mathbf{u}})$, some ${\mathbf{u}}\in\Sigma_{3}^{\omega}$ where $g=[a,b,c]$, some $a,b,c\in\Sigma_{3}^{+}$ such that:
@@
    Word $b$ is a prefix of $a$, which is a prefix of $c$.
@@
    Words $b$ and $c$ have a common suffix $s$ such that $|s|\geq|b|/2$.
Then a final segment of $\mathbf{w}$} has the form $g(\gamma^{2}({\mathbf{v}}))$, some ${\mathbf{v}}\in\Sigma_{3}^{\omega}$; equivalently, a final segment of $\mathbf{w}$} has the form $g^{\prime}({\mathbf{v}}))$ some ${\mathbf{v}}\in\Sigma_{3}^{\omega}$, where $g^{\prime}=c(g\circ\gamma^{2})c^{-1}=[cab,ca,caba]$. Letting $A=g^{\prime}(0)$, $B=g^{\prime}(1)$, and $C=g^{\prime}(2)$, we have
@@
    Word $B$ is a prefix of $A$, which is a prefix of $C$.
@@
    Words $B$ and $C$ have a common suffix $S$ such that $|S|\geq|B|/2$.
@@
Without loss of generality, replacing $\mathbf{w}$} by one of its final segments if necessary, suppose that $\mathbf{w}$} has the form $g({\mathbf{u}})$. By Lemma 6, words $00$ and $11$ are not factors of $\mathbf{u}$}. Therefore, if $2$ is not a factor of some final segment of $\mathbf{u}$}, then that final segment is $(01)^{\omega}$ or $(10)^{\omega}$. This is impossible since $\mathbf{w}$} is low. We conclude that every final segment of $\mathbf{u}$} contains a $2$.
Parse a final segment of $\mathbf{u}$} into blocks starting with $2$, and containing a single $2$. Words $21$ and $22$ are not factors of $\mathbf{u}$}; thus the blocks begin with $20$. Since $00$ and $11$ are not factors, 0’s and $1$’s alternate in the blocks. The block $20101$ is impossible; it would imply a block $z=2010101$ or $z=201012$. In both cases, $g(z)$ begins with cababacababa, since $a$ is a prefix of $c$. However, ababaababa is an $r$ power, $r\geq 5/2$. It follows that a final segment of $\mathbf{u}$} is concatenated from blocks $20$, $201$, and $2010$. Therefore, a final segment of $\mathbf{w}$} is concatenated from blocks $ca=A$, $cab=B$, and $caba=C$, and has the form $g^{\prime}({\mathbf{v}})$ for some ${\mathbf{v}}\in\Sigma_{3}^{\omega}$.
We see that property (1) certainly holds. Also,
@@
since $|c|\geq|b|$, establishing (2).
Finally, $C$ and $B$ have the common suffix $S=sa$. However,
@@
Define morphisms $g_{n}$ recursively by $g_{1}=g$, and $g_{n+1}=cg_{n}\circ\gamma^{2}c^{-1}$. By induction on the previous lemma, for each positive integer $n$, a final segment of $\mathbf{w}$} has the form $g_{n}({\mathbf{v}})$, some ${\mathbf{v}}\in\Sigma_{3}^{\omega}$.
Having a final segment of the form $g_{n+1}({\mathbf{v}})$ is equivalent to having a final segment of the form $g_{n}(\gamma^{2}({\mathbf{v}}))$. By induction on $n$, having a final segment of the form $g_{n}({\mathbf{v}})$ is equivalent to having a final segment of the form $g(\gamma^{2n}({\mathbf{v}}))$. Thus a final segment of $\mathbf{w}$} has the form $g(\gamma^{2n}({\mathbf{v}}))$. ∎
@@
Replacing $\mathbf{z}$} by a final segment if necessary, suppose that $\mathbf{z}$} has the form $h({\mathbf{u}})$. Let $S=\{z\in\Sigma_{3}^{*}:h(z)\text{ is a factor of }{\mathbf{z}}\}.$ Let $S^{R}=\{z^{R}:z\in S\}$. Then $S^{R}$ is an infinite factorial language. Using König’s Tree Lemma, choose a word ${\mathbf{z}}^{\prime}\in\Sigma_{3}^{\omega}$ such that every prefix of $\mathbf{z}$$^{\prime}$ is in $S^{R}$.
Define $h^{R}=[h(0)^{R},h(1)^{R},h(2)^{R}]$. Let ${\mathbf{Z}}=h^{R}({\mathbf{z}}^{\prime}).$ Then ${\mathbf{w}}={\mathbf{Z}}$ and $g=h^{R}$ satisfy the conditions of the Morphism Lemma, so that for every $g\in{\mathbf{G}}$, $h^{R}(g)$ is a factor of ${\mathbf{Z}}=h^{R}({\mathbf{z}}^{\prime})$.
Let $p$ be a prefix of $\mathbf{z}$$^{\prime}$ such that $h^{R}(g)$ is a factor of $h^{R}(p).$ Since every prefix of $\mathbf{z}$$^{\prime}$ is in $S^{R}$, write $p=z^{R}$ where $z\in S$. Then $h^{R}(g)$ is a factor of $h^{R}(z^{R})$. Thus $h(g^{R})$ is a factor of $h(z)$. By the definition of $S$, $h(z)$ is a factor of $\mathbf{z}$}, so that $h(g^{R})$ is a factor of $\mathbf{z}$}. ∎

### v3

@@ 5 Proof of the structure theorem
We prove the structure theorem via an intricate case analysis. Each case is determined by a set $S\subseteq F$, where $F$ is the set of ternary words which are not factors of $\mathbf{w}$}. As remarked in the previous section, it suffices to resolve the subcases $S=\{01,12,20,00\}$ and $S=\{11,22\}$.
In a given case, labelled by a set $S$, we proceed in one of three ways. We either
@@
Let a set $S$ of forbidden factors be given. Say that a word $w$ is good if it is low and has no factor in $S$. The concrete recipe for carrying out 1, 2, and 3 above is as follows.
@@
    Perform a backtrack search, looking for a good word $w$ of length 250. Word $w$ is a hypothetical prefix of $\mathbf{w}$}. If no such word $w$ exists, the case cannot arise.
@@
        If there are more than $2n+1$ needed factors then no low word of length 250 can be a prefix of $\mathbf{w}$}, since the condition on the complexity of $\mathbf{w}$} is violated. The case where $\mathbf{w}$} omits factors in $S$ cannot arise.
@@
    2. b.
      If no such $n$ is found, we look for a morphism satisfying the conditions of the morphism lemma or the dual morphism lemma. To do this
      1. i.
        Consider the good words $b$ with $|b|\leq 3$.
      2. ii.
        Suppose that for some positive integer $d\leq 200$, every good word of length $d+1$ has $b$ as a factor. This can be checked by a finite search.
      3. iii.
        Then consider the set $B=\{bu\in\Sigma_{3}^{*}:bub$ is good, $|u|\leq d$, and whenever $p$ is a prefix of bubbub such that $b$ is a suffix of $p$, then $p=b$ or p=bub}p=bub\}. By the first parsing lemma, $\mathbf{w}$} is concatenated from words of $B$.
      4. iv.
        We form a morphism using the blocks in B if either:
        • Set $B$ contains exactly 3 $b$-blocks;
        • Set $B$ contains exactly 4 $b$-blocks, including a $b$-block $s$ such that for all $r,t\in B$, word rstrst is not good. In this case we replace $B$ by $B-\{s\}$. By the second parsing lemma, a final segment of $\mathbf{w}$} is concatenated from the $b$-blocks of $B$.
      We form a morphism from the three $b$-blocks of $B$, and consider this morphism and its conjugates. If one of the considered morphisms satisfies the conditions of the morphism lemma or the dual morphism lemma the case is resolved.
@@
A python implementation of the above scheme in SageMath 9.3 (on a Microsoft Surface with Intel(R) Core(TM) i5-1035G4 CPU at 1.10GHz) resolved the case $S=\{11,22\}$ in about 36 seconds, and the case $S=\{01,12,20,00\}$ in about 9 seconds. The code is available at https://github.com/drjamescurrie-ai/RACE-3/tree/main. ∎
@@
In fact, the implementation mentioned above kept track of sets $S$ and the reasoning at each of its steps, and produced ‘human readable’ proofs. We give the reasoning that was output for the case $S=\{01,12,20,00\}$ in an appendix. Further, the implementation kept track of all the morphisms $h$ used. It may be the case that for some of these morphisms, the word $h({\mathbf{G}})$ is eligible and has critical exponent $\mu$.
@@ 6 Proofs of Lemmas
The critical exponent of $\mathbf{u}$} is irrational, and therefore is not realized by any of its finite factors. Thus for all $\epsilon_{2}>0$ and any positive integer $N$, the word $\mathbf{u}$} has a factor $uuu^{\prime}$ such that $u^{\prime}$ is a prefix of $u$ with $|u^{\prime}|>N$ and
@@
Proof of Lemma 3.
Let $\mathbf{s}$} be a suffix of $\mathbf{w}$}. It suffices to show that $\mathbf{s}$} contains letter 0.
Suppose $\mathbf{s}$} contains only letters 1 and 2. It must contain factors 12 and 21; otherwise $\mathbf{s}$} would end with $1^{\omega}$ or $2^{\omega}$ and therefore contain $5/2$ powers.
Suffix $\mathbf{s}$} must also contain factor 11 (and, symmetrically, factor 22): If not, assume $\mathbf{s}$} starts with 1, replacing $\mathbf{s}$} with one of its suffixes if necessary. Then $\mathbf{s}$} is concatenated from the $1$-blocks of $\mathbf{w}$}. Since 222 is a $3$ power, the only possible $1$-blocks of $\mathbf{w}$} are 12 and 122. The $1$-block 12 can only be used once in this concatenation, or else it appears in the context $2\underline{12}12$, which is a $5/2$ power. Then a suffix is $(122)^{\omega}$, which contains $5/2$ powers. This is a contradiction. Thus 11 must indeed occur.
So far we have shown that $\mathbf{s}$} contains factors 11, 12, 21, and 22. Since $\mathbf{w}$} contains $2(2)+1=5$ factors of length 2, there is exactly one length-2 factor of $\mathbf{w}$} containing a 0. It follows that the only 0 in $\mathbf{w}$} is its first letter. We thus see that $\mathbf{w}$} contains no factor of the form $a0$ where $a\in\Sigma_{3}$.
Consider the following set of 10 length-4 words:
$S=\{1121,1122,1211,1212,1221,2112,2121,2122,2211,2212\}.$
Let $s$ and $t$ be distinct words of $S$. Suppose $u$ is a low word over $\Sigma_{3}$ such that the set of factors of $u$ does not include any of 00, 10, 20, $s$, or $t$. For each of the $\binom{10}{2}$ possibilities for $s$ and $t$, a backtrack search shows that the length of $u$ is at most 39. It then follows that $\mathbf{w}$} contains 9 out of 10 words of $S$ as factors. In addition,the word $\mathbf{w}$} contains a length-4 factor starting with 0. However, the word $\mathbf{w}$} is eligible, and cannot contain $10$ length-4 factors. This is a contradiction. ∎
We use two preliminary lemmas in the proof of the morphism lemma (Lemma 4).
@@
Suppose that ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ is low. Let $a,b,c\in\Sigma_{3}^{+}$, and suppose that a suffix of $\mathbf{w}$} has the form $h({\mathbf{u}})$, some ${\mathbf{u}}\in\Sigma_{3}^{\omega}$, where $h=[a,b,c]$. Suppose further that
@@
    The word $b$ is a prefix of $a$, which is a prefix of $c$.
@@
    The words $b$ and $c$ have a common suffix $s$ such that $|s|\geq|b|/2$.
Then $\mathbf{u}$} doesn’t contain any of the factors $00$, $11$, 2121, and $22$.
@@
We treat the factors one at a time. For each factor we rule out 1-letter right extensions of that factor.
Factor $11$: If $11$ is a factor of $\mathbf{u}$}, then so is a factor $v$, where $v=110$, $v=111$, or $v=112$. In each case, $h(v)$ has the prefix bbbbbb which is a $3$ power. Since $\mathbf{w}$} doesn’t contain $3$ powers, the factor $11$ doesn’t occur in $\mathbf{u}$}.
Factor $00$: If $00$ is a factor of $\mathbf{u}$}, then so is a factor $v$, where $v=000$, $v=001$, or $v=002$. In each case $h(v)$ has the prefix aabaab since $b$ is a prefix of $a$ and $c$. However, aabaab is an $r$ power with $r\geq 5/2$, since $b$ is a prefix of $a$ and $|b|\geq|c|/2\geq|a|/2$.
Factor $21$: If $21$ is a factor of $\mathbf{u}$}, then so is a factor $v$, where $v=211$, $v=210$, or $v=212$. In each case, $h(v)$ has the prefix cbbcbb, since $b$ is a prefix of $a$ and $c$. This has the suffix sbbsbb, since $s$ is a suffix of $c$. But sbbsbb is an $r$ power with $r\geq 5/2$.
Factor $22$: If $22$ is a factor of $\mathbf{u}$}, then so is a factor $v$, where $v=220$, $v=221$, or $v=222$. In each case, $h(v)$ has the prefix ccbccb, which is an $r$ power with $r\geq 5/2$, since $b$ is a prefix of $c$ and $|b|\geq|c|/2$. ∎
It is convenient to work with $\gamma^{2}=[012,02,0102]$ rather than simply $\gamma$. If $h=[a,b,c]$, then $h\circ\gamma^{2}=[abc,ac,abac]$, and each image of a letter under $h\circ\gamma^{2}$ ends in $c$. This allows us to define $h^{\prime}=c(h\circ\gamma^{2})c^{-1}=[cab,ca,caba]$.
@@
Suppose that ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ is low. Suppose that $\mathbf{w}$} has the form $h({\mathbf{u}})$, some ${\mathbf{u}}\in\Sigma_{3}^{\omega}$ where $h=[a,b,c]$, some $a,b,c\in\Sigma_{3}^{+}$ such that
@@
    The word $b$ is a prefix of $a$, which is a prefix of $c$.
@@
    The words $b$ and $c$ have a common suffix $s$ such that $|s|\geq|b|/2$.
Then a suffix of $\mathbf{w}$} has the form $h(\gamma^{2}({\mathbf{v}}))$, some ${\mathbf{v}}\in\Sigma_{3}^{\omega}$; equivalently, a suffix of $\mathbf{w}$} has the form $h^{\prime}({\mathbf{v}}))$ some ${\mathbf{v}}\in\Sigma_{3}^{\omega}$, where $h^{\prime}=c(h\circ\gamma^{2})c^{-1}=[cab,ca,caba]$. Letting $A=h^{\prime}(0)$, $B=h^{\prime}(1)$, and $C=h^{\prime}(2)$, we have
@@
    The word $B$ is a prefix of $A$, which is a prefix of $C$.
@@
    The words $B$ and $C$ have a common suffix $S$ such that $|S|\geq|B|/2$.
@@
Replacing $\mathbf{w}$} by one of its suffixes if necessary, suppose that $\mathbf{w}$} has the form $h({\mathbf{u}})$. By Lemma 8, words $00$ and $11$ are not factors of $\mathbf{u}$}. Therefore, if $2$ is not a factor of some suffix of $\mathbf{u}$}, then that suffix is $(01)^{\omega}$ or $(10)^{\omega}$. This is impossible since $\mathbf{w}$} is low. We conclude that every suffix of $\mathbf{u}$} contains a $2$.
Parse a suffix of $\mathbf{u}$} into blocks starting with $2$, and containing a single $2$. Words $21$ and $22$ are not factors of $\mathbf{u}$}; thus the blocks begin with $20$. Since $00$ and $11$ are not factors, 0’s and $1$’s alternate in the blocks. The block $20101$ is impossible; it would imply a block $z=2010101$ or $z=201012$. In both cases, $h(z)$ begins with cababacababa, since $a$ is a prefix of $c$. However, ababaababa is an $r$ power with $r\geq 5/2$. It follows that a suffix of $\mathbf{u}$} is concatenated from blocks $20$, $201$, and $2010$. Therefore, a suffix of $\mathbf{w}$} is concatenated from blocks $ca=A$, $cab=B$, and $caba=C$, and has the form $h^{\prime}({\mathbf{v}})$ for some ${\mathbf{v}}\in\Sigma_{3}^{\omega}$.
We see that property 1 certainly holds. Also,
@@
since $|c|\geq|b|$, establishing 2.
Finally, the words $C$ and $B$ have the common suffix $S=sa$. However,
@@
Define morphisms $h_{n}$ recursively by $h_{1}=h$, and $h_{n+1}=ch_{n}\circ\gamma^{2}c^{-1}$. By induction on the previous lemma, for each positive integer $n$, a suffix of $\mathbf{w}$} has the form $h_{n}({\mathbf{v}})$, some ${\mathbf{v}}\in\Sigma_{3}^{\omega}$.
Having a suffix of the form $h_{n+1}({\mathbf{v}})$ is equivalent to having a suffix of the form $h_{n}(\gamma^{2}({\mathbf{v}}))$. By induction on $n$, having a suffix of the form $h_{n}({\mathbf{v}})$ is equivalent to having a suffix of the form $h(\gamma^{2n}({\mathbf{v}}))$. Thus a suffix of $\mathbf{w}$} has the form $h(\gamma^{2n}({\mathbf{v}}))$. It follows that whenever $g$ is a prefix of $\mathbf{G}$}, then $h(g)$ is a factor of $\mathbf{w}$}, and the lemma follows. ∎
@@
Replacing $\mathbf{z}$} by a suffix if necessary, suppose that $\mathbf{z}$} has the form $h({\mathbf{u}})$. Let $T=\{z\in\Sigma_{3}^{*}:h(z)\text{ is a factor of }{\mathbf{z}}\}.$ Let $T^{R}=\{z^{R}:z\in S\}$. Then $T^{R}$ is an right factorial language. Using König’s lemma, choose a word ${\mathbf{z}}^{\prime}\in\Sigma_{3}^{\omega}$ such that every prefix of $\mathbf{z}$$^{\prime}$ is in $T^{R}$.
Define $h^{R}=[h(0)^{R},h(1)^{R},h(2)^{R}]$. Let ${\mathbf{Z}}=h^{R}({\mathbf{z}}^{\prime}).$ Then ${\mathbf{w}}={\mathbf{Z}}$ and $h^{\prime}=h^{R}$ satisfy the conditions of the morphism lemma, so that for every $g\in{\mathbf{G}}$, word $h^{R}(g)$ is a factor of ${\mathbf{Z}}=h^{R}({\mathbf{z}}^{\prime})$.
Let $p$ be a prefix of $\mathbf{z}$$^{\prime}$ such that $h^{R}(g)$ is a factor of $h^{R}(p).$ Since every prefix of $\mathbf{z}$$^{\prime}$ is in $T^{R}$, write $p=z^{R}$ where $z\in T$. Then $h^{R}(g)$ is a factor of $h^{R}(z^{R})$. Thus $h(g^{R})$ is a factor of $h(z)$. By the definition of $T$, $h(z)$ is a factor of $\mathbf{z}$}, so that $h(g^{R})$ is a factor of $\mathbf{z}$}. ∎
Proof of first parsing lemma.
Let $\mathbf{w}$} be an eligible right-infinite word which omits the words of $S$. Every factor of $\mathbf{w}$} of length $d+1$ is low, and omits words of $S$. By the condition of the lemma it cannot omit $b$. Thus $b$ appears infinitely often in $\mathbf{w}$}, so that a suffix can be concatenated from the $b$-blocks of $\mathbf{w}$}. These blocks have the form $bu$ where
  • The word bubbub is a factor of $\mathbf{w}$};
Such a word $u$ is a factor of $\mathbf{w}$} and is therefore low and omits words of $S$. If $|u|>d$ it must therefore have $b$ as a factor. In this case, write $u=vbz$. Then $p=bvb$ is a prefix of bubbub distinct from $b$ and bubbub. This is a contradiction. ∎
Proof of second parsing lemma.
Write ${\mathbf{w}}=b_{1}b_{2}b_{3}b_{4}\cdots$ such that the $b_{i}\in B$. Then for $j\geq 2$, if $b_{j}=s$ let $b_{j-1}=r$ and $b_{j+1}=t$. Since rstrst is a factor of $\mathbf{w}$} it is low and cannot contain a factor from $S$. This is a contradiction. We conclude that the suffix $b_{2}b_{3}b_{4}\cdots$ of $\mathbf{w}$} is concatenated from word of $B-\{s\}$. ∎